OpenStudy (erinweeks):

Find the exact value of the trigonometric function. Do not use a calculator.

OpenStudy (erinweeks):

\[\cot (-\frac{ 5\pi }{ 4 })\]

OpenStudy (wolf1728):

180° = PI radians 90° = PI/2 radians and 45° = PI/4radians so you are asked to find the cot (-5 *45°) The cotangent of 45° = 1

OpenStudy (erinweeks):

thank you so much can i ask you one more quick question.

OpenStudy (wolf1728):

sure

OpenStudy (erinweeks):

The given angle is in standard position. Determine the quadrant in which the angle lies. -25° A. Quadrant III B. Quadrant II C. Quadrant IV D. Quadrant I

OpenStudy (wolf1728):

so you are asking about -25° and this has nothing to do with the -5*PI()/4 question right?

OpenStudy (erinweeks):

No its a whole different question... you weren't done with the 1st question?

OpenStudy (wolf1728):

I'd say it is asking for the cot (-5*45°) or cot (-225°) and that equals -1

OpenStudy (erinweeks):

i got the quadrant one already so dont mind the second question.. so the first question i asked is -1

OpenStudy (wolf1728):

As for the quadrant in which -25° is found I'd say quadrant IV (see attached)

OpenStudy (erinweeks):

yea i already got IV i figured that one out.. i appreciate the help anyways..

OpenStudy (wolf1728):

Yesd I'd day -1 is the answer to the first question. By the way did you find -25° is in quadrant IV? I guess you did !!

OpenStudy (erinweeks):

yea i found it thanks anyways!! and thank you for the first question as well!

OpenStudy (wolf1728):

you are welcome - and I hope I got it right (I think so - but you never know)